SENSITIVE FLUORESCENCE SPECTROSCOPY OF JET COOLED ^{15}NO_{2}

E. A. VOLKERS, A. VREDENBORG, \underline{H. LINNARTZ}, J. BULTHUIS, S. STOLTE, Laser Centre Vrije Universiteit, Department of Physical Chemistry, De Boelelaan 1083, NL 1081 HV Amsterdam, The Netherlands; R. JOST, Laboratoire de Spectrometrie Physique, Universite Joseph Fourier de Grenoble, BP 87, 38402 Saint Martin d’Heres Cedex, France.

A spectroscopic setup designed for high resolution spectroscopy of jet cooled ^{x}O^{y}N^{z}O isotopologues is described with the aim to obtain high quality laboratory spectra for the study of the mass independent effect in this triatomic molecule. A special piezo valve allows operation at minimal gas consumption and time gated fluorescence spectroscopy is used as a highly sensitive detection technique. The performance of the setup is demonstrated on the first A^{2}B_{2} - X^{2}A_{1} rovibronic transitions of ^{15}N^{16}O_{2} measured in a jet. The gas consumption is as low as 0.025 mg per cm^{−1} spectral range. Special efforts have been made to extend applications to lower energies, using a LN_{2} cooled germanium detector.

[1] E.A. Volkers, A. Vredenborg, H. Linnartz, J. Bulthuis, S. Stolte, R. Jost, Chem. Phys. Lett., submitted (2004).